OpenStudy (mtalhahassan2):

Need help!! 2. Give that cos (x+y) = cos x cos y – sin x sin y, use this to express cos 2x in terms of cos x

OpenStudy (campbell_st):

so replace y with x so you have cos(x + x) = cos(x)cos(x) - sin(x)sin(x) and simplify from there.

OpenStudy (campbell_st):

so the left side of the equation becomes cos(2x)

OpenStudy (mtalhahassan2):

why do you do that

OpenStudy (campbell_st):

now you can simplify the right side

OpenStudy (mtalhahassan2):

I mean why do you change the y into x

OpenStudy (campbell_st):

that way when you replace y with x cos(x + y) = cos(x + x) which is cos(2x)

OpenStudy (mtalhahassan2):